OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). Answer y = 4/3x + 8 y = 4/3 x + 3 y = -3/4 x + 8 y = -3/4 x + 3

OpenStudy (anonymous):

\[y=4/3x+8\]

OpenStudy (callisto):

// to the line => same slope Slope of line 4x - 3y = -12 = -(4/-3) = 4/3 Equation required is y-4 = (4/3) [x-(-3)] y-4 = (4/3)x + 4 y = (4/3)x + 4+4 y = (4/3)x + 8